OpenStudy (anonymous):

32sqrt(x*x^2*x^5)???

OpenStudy (anonymous):

HELP PLEASE?!?!?

OpenStudy (anonymous):

@becca_ashley12

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ a^n\cdot a^m\implies a^{n+m}}}\\ \quad \\ \quad \\ 32\sqrt{x\cdot x^2\cdot x^5}\implies 32\sqrt{x^1\cdot x^2\cdot x^5}\implies 32\sqrt{x^{1+2+5}}\)

OpenStudy (anonymous):

so... um x^8 is under the radicand bar, right?

OpenStudy (jdoe0001):

recall the exponents rule, same base, different exponents, add the exponents, keep the base -> http://www.math-play.com/image-exponents-rules.jpg yes, will be \(x^8\)

OpenStudy (anonymous):

how do I do 32sqrt(x^8)?

OpenStudy (jdoe0001):

then take out what you can from the radicand :)

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ a^n\cdot a^m\implies a^{n+m}}}\\ \quad \\ \quad \\ 32\sqrt{x\cdot x^2\cdot x^5}\implies 32\sqrt{x^1\cdot x^2\cdot x^5}\implies 32\sqrt{x^{1+2+5}}\\ \quad \\ 32\sqrt{x^8}\implies \large 32\sqrt[{\color{red}{ 2}}]{(x^4)^{\color{red}{ 2}}}\)

OpenStudy (anonymous):

so... um 32x^4???

OpenStudy (jdoe0001):

so... um 32x^4??? \(\large \checkmark\)

OpenStudy (anonymous):

ok, Thanks sooo much!!!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

:)